# SVLA: A Unified Speech-Vision-Language Assistant with Multimodal Reasoning and Speech Generation

### v1

@@ 4.4 Speech Decoding
Table 2: Speech-Based Captioning and VQA Performance Comparison
Model Input $\rightarrow$ Ouput Image Captioning (Similarity) VQA (Speech-VQA-Accuracy)
COCO-2014 COCO-2017 Flickr30k VQAv2-val OKVQA-test GQA-test VizWiz
[…]
4.5 Text-Instruction Outputs
During the pre-train stages (can be found in Appendix A.2), the model generate text tokens and speech tokens directly. Similarly, during the SFT stage, the model generates responses directly in textual form. However, for speech output, we use a structured prompting system to guide the model in reasoning and generating coherent spoken responses. Unlike text generation, speech output introduces additional complexities, requiring structured guidance to ensure fluency, coherence, and semantic alignment. For ASR and TTS tasks, we use simple yet effective prompts to map between speech and text. ASR follows “The transcript of the given speech is:” to extract text, while TTS uses “This is how your text sounds in speech:” to generate speech. For speech-vision tasks, where the model must generate speech from visual inputs, we adopt an instruction-based speech generation approach. Here, textual captions and textual answers serve as intermediate reasoning steps to align the model’s understanding before producing spoken outputs. In image captioning, we explicitly structure the prompt as “The textual caption is ‘caption’. Therefore, the audio caption is:” followed by discrete speech tokens. This ensures that the speech output remains faithful to the structured textual representation. Similarly, for VQA tasks, we prompt the model with “The textual answer is ‘answer’. Therefore, the audio answer is:” before generating discrete speech tokens. This method enforces a structured reasoning process, ensuring that the model first forms a grounded textual response before converting it into speech.
5 Experiments
5.1 Implementation Details
We use Qwen2.5-1.5B as our backbone LLM and train it with PyTorch. The vision encoder is CLIP-large-patch14-336 [44], which producing $256$ total tokens for each image. Speech data is handled by the SpeechTokenizer[43], which encodes each 1-second segment of audio into 50 discrete tokens; when advanced speech fidelity is required, SoundStorm-SpeechTokenizer [^2] https://github.com/ZhangXInFD/soundstorm-speechtokenizerextends the same quantization principle to more nuanced tasks. The weights of these speech components are from AnyGPT[12]. More complementary details can be found in Appendix B.
5.2 Metrics
Text-Output tasks:
Speech-output tasks:
We evaluate speech generation using two methods: We transcribe model-generated speech with Whisper Medium [48] and compare the resulting text to human references (as in text-output tasks). However, ASR models can exhibit biases that introduce transcription errors, thereby distorting the perceived quality of the generated speech. Therefore, we use WavLM-TDNN [41] to extract speech embeddings from both generated and reference speech and measure their similarity (e.g., via cosine similarity). This directly compares acoustic properties without relying on ASR. Apart from that, we propose a Speech-Vqa-Accuracy that incorporates the above similarity. Given a generated answer $\hat{A}$ and $N$ human-annotated references $A_{1},A_{2},\dots,A_{N}$, let the set of unique references be $\{U_{1},U_{2},\dots,U_{n}\}$ (where $n\leq N$). We compute the similarity $\mathrm{sim}(\hat{A},U_{j})$ for each $U_{j}$, and choose the closest reference:
$U_{\mathrm{best}}=\arg\max_{U_{j}}\,\mathrm{sim}(\hat{A},U_{j}).$ (3)
Let $\mathrm{count}(U_{\mathrm{best}})$ be the number of times $U_{\mathrm{best}}$ appears in the original $N$ references. We define a weight:
$W(U_{\mathrm{best}})=\begin{cases}1,&\text{if }\mathrm{count}(U_{\mathrm{best}})\geq 0.3N,\\ \tfrac{2}{3},&\text{if }\mathrm{count}(U_{\mathrm{best}})=0.2N,\\ \tfrac{1}{3},&\text{if }\mathrm{count}(U_{\mathrm{best}})=0.1N.\end{cases}$ (4)
Finally, we define the Speech-VQA Score:
$\mathrm{Score}_{\mathrm{VQA}}=I\!\bigl{(}\mathrm{sim}(\hat{A},U_{\mathrm{best}})\,\geq\,\tau\bigr{)}\,\cdot\,W(U_{\mathrm{best}}),$ (5)
where $\tau$ is a similarity threshold, and $I(\cdot)$ is an indicator function that is 1 if the condition is true, and 0 otherwise. In this paper, we set $\tau$ to 0.7 based on empirical observations that this threshold best balances precision and recall, effectively distinguishing perceptually similar speech segments from dissimilar ones.
5.3 Results
Figure 3: Visualization of the model’s attention maps when aligning image features with text (and/or spoken) queries.
Model Backbone ASR TTS
LibriSpeech VCTK
WER WER Similarity
Human-level - 5.8 1.9 0.93
Wav2vec 2.0 [38] - 2.7 - -
Whisper Large V2 [48] - 2.7 - -
VALL-E [50] - - 7.9 0.75
VILA [51] Mixtral 8x7B [52] 8.1 - -
AnyGPT-7B LLaMA-2-7B 8.5 8.5 0.77
MIO-Ins [47] Yi-6B-Base [53] 10.3 4.2 -
SVLA-2B Qwen-1.5B 10.2 21.7 0.69
SVLA-2B-Text-Ins Qwen-1.5B 9.2 11.2 0.74
Text-Speech Performance:
From Table 3, SVLA-2B-Text-Ins outperforms its direct-speech counterpart, SVLA-2B, on both ASR and TTS tasks. By prompting the model with structured text (as “This is the transcript:” and “This is how the text sounds:”), it reduces ASR WER from 10.2 to 9.2 and TTS WER from 21.7 to 11.2, and increases similarity from 0.63 to 0.74. Although both models still lag behind specialized ASR systems like Whisper Large V2 and Wav2vec 2.0 (each with WER 2.7), SVLA-2B-Text-Ins is competitive with the larger AnyGPT-8B (ASR WER 8.5, TTS WER 8.5, similarity 0.77). Direct speech generation (SVLA-2B) avoids verbose prompts but can sacrifice clarity, whereas instruction-based speech generation (SVLA-2B-Text-Ins) leverages explicit textual guides to improve transcription and synthesis accuracy, demonstrating the trade-off between naturalness and performance.
Image Captioning Performance:
From Table 1, instruction tuning in SVLA-2B-Text-Ins has minimal impact on text-only performance, evidenced by near-identical CIDEr scores on COCO-2014 (120.0 vs. 120.2) relative to the non-instruction version, SVLA-2B. However, when generating spoken captions, SVLA-2B scores close to zero, whereas SVLA-2B-Text-Ins jumps to around 56—an improvement driven by structured prompts (as “The textual caption is … Therefore, the audio caption is: …”) that significantly boost speech accuracy. This advantage is also clear in acoustic similarity metrics: on COCO-2014, SVLA-2B achieves 79.7, while SVLA-2B-Text-Ins rises to 89.3, reflecting more faithful speech outputs (Table 2). While TMT attains an even higher speech-only result (78.7 of CIDEr scores on COCO) by treating each modality as a separate “language”, it handles only one input modality at once and thus lacks a unified approach to image+text or image+speech inputs. Moreover, Next-GPT-7B surpasses SVLA-2B in text tasks (158.3 vs.  120 CIDEr) but does so at a larger 7B scale, while SVLA-2B (Qwen2.5-1.5B) maintains parameter efficiency and handles both text and speech.
VQA Performance:
In Table 1 shows that SVLA-2B-Text-Ins stands out in the I+T → T VQA task by delivering competitive performance despite utilizing a much smaller Qwen-1.5B backbone compared to the 7B-sized alternatives. While models like LLaVA-1.5 and Next-GPT-7B achieve strong scores on benchmarks such as VQAv2, OKVQA, and GQA, our SVLA-2B-Text-Ins variant demonstrates impressive results with a VQAv2 score of 69.7 and a VizWiz accuracy of 58.0, rivaling or even surpassing larger models. On the other hand, the model shows a significant drop—by approximately 20% in OKVQA and GQA—when evaluated on the I+S → T VQA task, indicating that the model performs better when the question is provided in text form. Notably, similar to the image captioning task, the instructive output strategy does not substantially impact performance in text-output VQA tasks, as the accuracy difference between non-instruct and instruct configurations is minimal.
The performance of speech outputs is generally worse than that of text outputs by roughly 15–20% on VQA tasks. In the ASR-based evaluation (Table1), the SVLA-2B-Text-Ins model operating in the I+T → S configuration achieves a VQAv2 score of 37% and an OKVQA score of 11.6% , compared to 29.4% and 6.08% , respectively, in the I+S → S configuration. This discrepancy indicates that integrating a text input—as opposed to directly processing speech—preserves richer semantic details. The acoustic Speech-VQA accuracy evaluation further highlights this benefit, with the SVLA-2B-Text-Ins model recording a VQAv2 accuracy of 51.7% and an OKVQA accuracy of 69.6% in I+T → S, versus 44.5% and 62.5% in I+S → S. These improvements are likely due to the structured intermediate text-generation stage, which reduces the ambiguity inherent in raw speech processing and results in more semantically faithful responses for complex VQA tasks.
5.4 Ablation
Where does the model look to the image when writing and speaking:
To investigate how the model allocates visual attention during text and speech generation, we visualize the attention patches corresponding to specific words in both modalities in Figure 3. In these examples, the text yields more focused attention than the speech. In Case A, text attention precisely highlights both individuals for “people” and localizes the “frisbee”, while speech attention is more diffused. In Case B, text cleanly isolates the rider for “man” and the wheels for “skateboard”, whereas speech spreads across a wider area. In Case C, text remains sharply concentrated on the rider for “man” and the animal for “horse”, but speech again shows broader contextual focus. This precision in text helps achieve better performance in visual tasks, as it zeroes in on the exact objects needed for accurate descriptions.
Effect of Similarity Threshold on Speech-based VQA
Figure 4: Impact of Similarity Threshold on acoustic Speech-VQA Accuracy.
@@ 5.4 Ablation
In Figure 4, we examine how different similarity thresholds (0.6 to 0.9) affect speech-based VQA accuracy. As the threshold increases, all models exhibit a sharp decline in accuracy, underscoring the sensitivity of speech-based metrics to acoustic matching. Notably, instruction-tuned models consistently outperform their non-instruction-tuned counterparts across all thresholds, suggesting that structured guidance improves both the semantic and acoustic quality of the generated speech. This result also highlights the importance of threshold selection: a higher threshold imposes stricter acoustic requirements, potentially penalizing answers that are semantically correct yet acoustically divergent from the references.
Does Text Guidance Make Speech Smarter?
According to Figure 5, SVLA-2B-Text-Ins generates more structured and coherent speech than SVLA-2B, enhancing semantic alignment and reasoning. By first generating a structured textual response, the model ensures that speech output follows a well-defined linguistic structure. This intermediate step helps maintain consistency, reduces ambiguity, and improves multimodal alignment, leading to clearer and more contextually relevant spoken responses. In contrast, SVLA-2B, which generates speech directly, tends to produce more random and less structured outputs due to the absence of textual guidance. This approach increases the likelihood of semantic misinterpretation, prosody misalignment, and overall incoherence.
@@
ASR-based evaluation falls short in assessing speech quality because it misinterprets minor phonetic variations as errors, even when the spoken response is nearly correct. As seen in Figure 5(a), the model’s speech output closely follows the intended text, yet ASR misrecognizes words like “camera” as “camaral” and “kimeroh” or “tan” as “tanh”. These small differences, though perceptually acceptable, lead to unfair penalization since ASR focuses on exact text matching rather than acoustic similarity. This limitation reveals the need for an evaluation method that directly measures speech waveform similarity rather than relying on error-prone transcription.
@@ 6 Conclusion
In this work, we introduce SVLA, a unified Speech-Text-Vision Assistant capable of handling both language tasks (ASR, TTS) and vision-language reasoning tasks (image captioning, VQA). To support the community in building similar models, we also release a large-scale tri-modal dataset encompassing speech, text, and vision. Additionally, we analyze two settings for speech generation: directly producing spoken output and using a text prompt to guide speech synthesis. We further propose Speech-VQA-Accuracy, a new metric based on acoustic similarity for evaluating spoken VQA responses. Our experiments show that while the model performs better with text outputs, speech outputs benefit from an instructive text prompt, yielding more coherent.
Our study has several limitations. First, the speech data in our dataset is generated using a TTS model, which may lack the natural emotional and prosodic variations found in real human speech. Second, our architecture relies on a straightforward concatenation-based fusion, whereas more sophisticated approaches—such as Q-Former [7] or Perceiver Resampler [9]—could potentially yield better performance. Finally, due to resource constraints, we build on Qwen2.5-1.5B, which limits the model’s capacity. We hypothesize that employing a larger backbone would further enhance speech-related capabilities.

### v2

@@ 5.3 Results
Image Captioning Performance: From Table 2, instruction tuning in SVLA-2B-Text-Ins leads to minimal change in text-only captioning performance. For instance, on COCO-2014, the CIDEr score improves only slightly from 120.0 (SVLA-2B) to 120.2. However, in the speech captioning setting (I+S→S), instruction tuning results in a substantial gain: SVLA-2B scores only 2.0, while SVLA-2B-Text-Ins reaches 62.2. This highlights the effectiveness of structured prompts (e.g., “The textual caption is … Therefore, the audio caption is:”) in guiding coherent speech generation. TMT [35], which performs I→S generation directly, achieves a CIDEr score of 78.7 on COCO-2014. While upper SVLA-2B-Text-Ins, it operates under a different paradigm—treating each modality independently—whereas SVLA supports unified multimodal reasoning across both text and speech outputs.
VQA Performance: Table 2 shows that SVLA-2B-Text-Ins performs competitively in the I+T→T VQA setting, despite using the smaller Qwen-1.5B backbone compared to 7B-scale models. While models like LLaVA-1.5 and Next-GPT-7B achieve strong performance on benchmarks such as VQAv2, OKVQA, and GQA, SVLA-2B-Text-Ins achieves comparable results, with a VQAv2 score of 69.7 and a VizWiz score of 58.0. These results demonstrate that a smaller, instruction-tuned model can rival or even surpass larger alternatives.
However, performance declines in the I+S→T setting, where the question is spoken. For example, SVLA-2B-Text-Ins drops to 52.7 on VQAv2 and 28.7 on OKVQA—approximately 15–20 points lower than in the I+T→T setting. This suggests that the model performs more effectively when the input question is provided in text rather than speech. Speech output accuracy is generally lower than text output across VQA tasks. In the I+T→S setting, SVLA-2B-Text-Ins achieves 37.5 on VQAv2 and 11.6 on OKVQA, whereas in I+S→S, the scores fall to 29.4 and 6.08, respectively. This drop highlights the difficulty of reasoning directly from speech inputs and generating accurate spoken responses. These results suggest that using text as an intermediate representation enhances semantic alignment, particularly for complex reasoning tasks. By contrast, SVLA-2B completely fails to handle VQA tasks.
5.4 Ablation Studies
Effect of Accent and Speaking Speed:
(a) I+T→S
(b) I+S→S
Figure 3: Impact of accent and speaking speed on VQA performance. Both plots show accuracy on VQAv2-val.
We evaluate the impact of different accents and speaking speeds on VQA accuracy using the SVLA-2B-Text-Ins model on the VQAv2-val set. As show in table 3, in both the I+S→T and I+S→S settings, American and British accents yield higher performance, while Indian and Australian accents result in noticeably lower accuracy. Notably, despite being included in the training data, the Indian accent still underperforms, indicating potential challenges in generalization or speech variability. In terms of speaking speed, the model performs best around the default rate (1.0×), with accuracy dropping at both extremes. The decline is most pronounced at 1.3×, suggesting that faster speech reduces recognition and reasoning quality.
Figure 4: Visualization of attention maps comparing SVLA-2B’s visual grounding accuracy with and without intermediate textual instructions during speech generation.
Where Do the Models Look in Images?:
To examine how different speech generation strategies affect visual grounding, we visualize the model’s attention maps in Figure 4. For each key word, we show three attention maps: the first green map is from the text output of SVLA-2B-Text-Ins, the second green map from its instruction-based speech output; and the blue map from SVLA-2B’s direct speech output.The text output from SVLA-2B-Text-Ins exhibits the most focused and accurate attention, precisely grounding visual entities. Its speech output (green) generally retains meaningful grounding, showing acceptable attention consistency. In contrast, the speech output from SVLA-2B (blue) is often unfocused or misaligned, failing to attend to the relevant image regions. These results demonstrate that textual instructions play a critical role in guiding the model’s visual attention. Without the intermediate text step, the model lacks semantic anchoring and often fails to locate the correct objects in the image, leading to degraded visual grounding during speech generation.
@@ 5.4 Ablation Studies
ASR-based evaluation falls short in assessing speech output quality, as it often misinterprets minor phonetic variations as errors—even when the spoken response is semantically accurate. As shown in Figure 5(a), the model’s speech output closely matches the intended text, yet ASR transcribes words like “camera” as “camaral”, or “tan” as “tanh”. These subtle differences, while perceptually acceptable, are unfairly penalized because ASR prioritizes exact word-level matching over acoustic or semantic similarity. This limitation underscores the need for more robust evaluation metrics that go beyond transcription accuracy. In particular, speech-based VQA and captioning tasks would benefit from metrics that directly assess the fidelity of the generated speech waveform—capturing both semantic correctness and acoustic naturalness—without relying solely on error-prone intermediate transcriptions.
6 Ethical Considerations
We utilize publicly available datasets containing licensed image-text and speech-text pairs. All speech samples are either synthetic or derived from open corpora, explicitly excluding personal or sensitive data. While our datasets incorporate diverse accents and varying speaking rates to enhance representativeness, synthetic speech may still not capture the complete variability inherent in natural human speech. Observed performance disparities across different conditions highlight the necessity of ongoing research in fairness and robustness.
7 Limitations:
Our study has several limitations. First, the speech data is generated using a TTS model, which may lack the natural prosody and emotional variation of real human speech. Despite augmentations in accent, speed, and noise, the resulting speech may still be less diverse than natural input. Second, we use the Qwen2.5-1.5B backbone due to resource constraints, which limits model capacity. Third, the speech tokenizer introduces decoding errors, reducing intelligibility even when the underlying text is accurate.
@@ 8 Conclusion
In this work, we introduce SVLA, a unified Speech-Text-Vision Assistant capable of handling both language tasks (ASR, TTS) and vision-language reasoning tasks (image captioning, VQA). To support the community in building similar models, we also release a large-scale tri-modal dataset encompassing speech, text, and vision. Additionally, we analyze two settings for speech generation: directly producing spoken output and using a text prompt to guide speech synthesis. Our experiments show that while the model performs better with text outputs, speech outputs benefit from an instructive text prompt, yielding more coherent. In future work, we plan to incorporate real human speech, improve speech tokenization quality, and explore larger model backbones to better support nuanced prosody, robustness to speech variability, and high-fidelity speech generation.